OpenStudy (anonymous):

The probability of winning a particular game at the Game Place is 1/4 . Find the odds against winning the game. A. 1 to 4 B. 3 to 4 C. 3 to 1 D. 5 to 1

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

Odd = 1 - P(win) = 1 - 1/4 = 3/4